Lynn and Laws knew that the confection should be exactly 8% sugar. It was to be made using one component that was 5% sugar and another that was 20% sugar. How much of each should they use to get 800 ml of the confection?
Answer by Nate(3500)

You should take the average:
( 5% * x + 20% * y ) / ( x + y ) = 8%
* x + y = 800
( 0.05x + 0.2y ) / 800 = 0.08
* x = 800 - y
( 0.05( 800 - y ) + 0.2y ) / 800 = 0.08
( 40 - 0.05y + 0.2y ) / 800 = 0.08
( 40 + 0.15y ) / 800 = 0.08
40 + 0.15y = 64
0.15y = 24
y = 160
Substitute: y = 160
x = 800 - y
x = 800 - 160 = 640
640ml of the 5% and 160ml of the 20%